45.

45 Problem 16-140 Tobac Company reported an operating loss of S139000 for financial report ro and tax purposes n ore he enacted ax rae is 40% for 2018 and all future years. Assume that Tobac elects a loss carryback. No valuation allowance is needed for any deferred tax assets. incone rates paid $11,lee 30% $12,see 35% $17,se 4$18,see 30% 2814 2015 2016 2017 37,eee $42,eee s49,000 $47,eee Print References Required: 1:-Prepare a compound journal entry to record Tobacs tax provision for the year 2018. 2 Compute Tobacs net loss for 2018. Complete this question by entering your answers in the tabs below Required 1 Required 2 Prepare a compound journal entry to record Tobacs tax provision for the year 2018. (If no entry is required for a transaction/event, select No journal entry raquired in the first account field.) Journal entry worksheet Record the income taxes <Pray 450, 50 İİİ Next >

req 1

Complete this question by entering your answers in the tabs below. Required 1 Required 2 eBook Prepare a cormpound journal entry to record Tobacs tax provision for the year 2018. (If no entry is required for a transaction/event, select No journal entry required in the first account field.) Print View transaction list Journal entry worksheet Record the income taxes Note: Enter debits before credits Transaction General Journal Debit Credit Receivable-Income tax refund Deferred tax asset ncome tax benefit

req 2

lobac Company reported an operating loss ot st y 000 tor tinancial reporting and tax purposes in 2018. I he enacted tax rate is 40% for 2018 and all future years. Assume that Tobac elects a loss carryback. No valuation allowance is needed for any deferred tax assets. Taxable income, tax rates, and income taxes paid in Tobacs first four years of operations were as follows: paid $11,100 38% 2014 $37,000 2015 $42,000 30% $12,600 2016 $49,000 35% $17,150 2017 $47,000 40% $18,800 eBook Print Required 1. Prepare a compound journal entry to record Tobacs tax provision for the year 2018. 2. Compute Tobacs net loss for 2018. Complete this question by entering your answers in the tabs below Required 1 Required 2 Compute Tobacs net loss for 2018. loss Required 1
Answer #1
1 Transaction General Journal Debit Credit
1 Receivable-Income tax refund (Note:1) 35950
Deferred tax asset (Note:2) 17200
Income tax benefit 53150
(Recorded tax provisions for 2018)
Note:1
Carryback loss of 2018 to 2016 and 2017
Year Amount Tax rate Tax
Refund
2016 49000 35% 17150
2017 47000 40% 18800
Total 35950
Carry forward balance loss of 2018 to 2019
Balance loss to be carry forward=139000-49000-47000=$ 43000
Deferred tax asset @ 40%=43000*40%=$ 17200
2 Net loss for 2018:
$
Loss for 2018 139000
Less: Income tax benefit 35950
Net loss 103050